This week marked the end of another dark run and thus the end of another science period. We are now also three quarters of the way through the current science period labeled 17-1. As such we are starting to see a larger number of filler Priority 4 targets being observed. In this last dark run we observed 374 acceptable observations for a total of 38 hours and 122 Priority 4 targets for a total of 22 hours. We saw 6% of the time lost due to problems. This is a fairly respectable number and fairly typical of the mature HET before the takedown.
We are now three quarters of the way through the current 17-1 period. The Telescope Allocation Committees have begun meeting and the new allocations are being submitted for the 17-2 period which begins April 1. For this period the same suite of instruments will be available for the PIs: LRS2-B and LRS2-R and the two weeks around dark time will also be fully committed to science observing.